# ECATS: Explainable-by-design concept-based anomaly detection for time series

### v1

@@ Abstract
Deep learning methods for time series have already reached excellent performances in both prediction and classification tasks, including anomaly detection. However, the complexity inherent in Cyber Physical Systems (CPS) creates a challenge when it comes to explainability methods. To overcome this inherent lack of interpretability, we propose ECATS, a concept-based neuro-symbolic architecture where concepts are represented as Signal Temporal Logic (STL) formulae. Leveraging kernel-based methods for STL, concept embeddings are learnt in an unsupervised manner through a cross-attention mechanism. The network makes class predictions through these concept embeddings, allowing for a meaningful explanation to be naturally extracted for each input. Our preliminary experiments with a simple CPS-based dataset show that our model is able to achieve great classification performance while ensuring local interpretability.
@@ 3 ECATS architecture
Here we describe ECATS (Explainable-by-design Concept-based Anomaly detection model for Time Series), the first interpretable neuro-symbolic concept-based model for trajectories. The goal of our proposed methodology is to discriminate between normal and anomalous signals, contextually providing an explanation for the outcome. In particular, the form of the explanation is a disjunction of interpretable STL formulae, which act as concepts in our model, whose semantics is influential for determining the label of the input trajectory. Indeed, as mentioned in Section 2.0.2, STL is a concise yet rich language to describe the characteristics of time series, and therefore is a valuable candidate to serve as a concept language in this field. In fact, small STL formulae describe simple patterns in a signal, and are easy to understand for a human (possibly once converted into a natural language description).
We now describe at a high level how the model works. We provide Figure 1 as a reference to graphically follow the discussion. Our model takes as input a signal $\xi$ and, through a mechanism of cross-attention (Subsection 3.1) between a kernel embedding of the concepts (computed as per Equation 1) and the input trajectory, creates an efficient representation of the relationship between the two. The end product gets fed to an MLP classifier net, which outputs its prediction $y$, the class of the input trajectory. Our model is then able to provide not only excellent classification accuracy (Section 4), but also easy-to-read explanations made of simple STL formulae (Subsection 3.2).
@@ 4.0.1 Setup
We evaluate ECATS on a train cruise control dataset proposed by [22], and use the standard training/test split. The dataset, balanced between classes, collects $200$ trajectories of $1$ dimension and has a clear-cut condition to classify between regular and anomalous trajectories, except for seven outliers. In Figure 4, a visual representation of the dataset.
Figure 2: Representation of the classes in the train cruise control dataset.
Figure 3: Robustness of global explanation for anomalous trajectories’ class.
Figure 4: An example of regular trajectory.
@@ 4.0.2 Evaluation
To assess the classification performance of our model, we run the tests using $5$ different initialisation seeds. This shows that its classification accuracy averages at $96.5\%$, with a standard deviation of $2.55$. Following, some example outputs from ECATS, starting from a local explanation. Consider the post-processed explanation given for a signal $\xi$ correctly classified as regular:
@@
Here is an example of post-processed global explanations given by ECATS:
@@
In Figure 4, the robustness values for the formula given as explanation for the anomalous class, computed for all trajectories. Except for some points which are most likely the aforementioned outliers, the class separation is linear. For training details and further discussion of results, we refer to Appendix E.
@@ Appendix A Signal Temporal Logic
Signal Temporal Logic (STL) [19] is a formal specification language to express temporal properties over real-value trajectories with a dense time interval. We now present its syntax and semantics.
@@ Appendix C Selecting the concepts
Once all the templates have been generated, we instantiate each of them on a set of parameter configurations $\mathcal{P}=\{p_{0},\ldots p_{|\mathcal{P}|}\}$ to be filtered by Signature-based Optimisation as defined in [20]. Denoting as $\varphi(p)$ the application of the parameter set $p$ to the template formula $\varphi$, and given a set of trajectories $\hat{\mathcal{T}}=\{\xi_{0},\ldots,\xi_{|\hat{\mathcal{T}}|}\}$ (with $p\in\mathcal{P}$, and $\hat{\mathcal{T}}\subset\mathcal{T}$), it consists in building a matrix $S\in\mathbb{R}^{|\mathcal{P}|\times|\hat{\mathcal{T}}|}$, whose rows are called the signature of the formula, s.t. $S(i,j)=\rho(\varphi(p_{i}),\xi_{j})$, i.e. the cell $(i,j)$ of such matrix holds the robustness of the $i$-th instantiation of $\varphi$ on the $j$-th trajectory. In such matrix, the distance between rows can be considered as a proxy for semantic similarity among different parameterisations of the same template, hence a selection on formulae can be done by keeping only those whose signature has a distance higher than a certain threshold $\tau\in\mathbb{R}_{\geq 0}$ from formulae already selected. This helps in removing redundancy in the set of concepts. In our case, we use signature-based optimisation with cosine distance and threshold $\tau=0.9$ and, for interpretability, we keep $M$ fairly small, namely $M=3$.
@@ Appendix E Training setup and results
First, we better analyse the local explanation for a regular trajectory outlined in Section 4 by reporting the plot of the robustness values for the considered signal versus all the signals of the opposite class (Figure 5). For ease, we report here the provided explanation:
$\neg\,(\xi\leq 35.9\ \mathcal{U}_{[11,36]}\ \xi\geq 31.5)$
Figure 5: Robustness values for the explanation of the regular trajectory reported as example in Section 4.
A per Figure 5, the provided formula does an excellent job at separating the signal belonging to the regular class from the ones that are anomalous. Coincidentally the cluster of values regarding the anomalous class has opposite sign with respect to the robustness on the considered regular signal, therefore the formula has no need to be shifted to increase human readability.
Figure 6: An example of anomalous trajectory.
$\mathcal{G}_{[0,24]}\ \xi\leq 37.3$
$\neg\,\mathcal{G}_{[0,24]}\ \xi\leq 28.7$
See Appendix D for more details on the simplification of formulae.
We now recall the global explanation by class reported in Section 4:
Regular: $\neg\,(\xi\geq 20.0\ \mathcal{U}_{[23,47]}\ \xi\leq 40.5)$
Anomaly: $\neg\,\mathcal{G}_{[0,24]}\ \xi\leq 28.7$
and report here the plot for the robustness values of the explanation for the regular class (Figure 8), in order to integrate the one already reported in Section 4, Figure 4. As mentioned earlier, Except for a few points which are most likely the outliers, the separation between classes is clear.
Figure 8: Robustness of global explanation for regular trajectories’ class.

### v2

@@ Abstract
Deep learning methods for time series have already reached excellent performances in both prediction and classification tasks, including anomaly detection. However, the complexity inherent in Cyber Physical Systems (CPS) creates a challenge when it comes to explainability methods. To overcome this inherent lack of interpretability, we propose ECATS, a concept-based neuro-symbolic architecture where concepts are represented as Signal Temporal Logic (STL) formulae. Leveraging kernel-based methods for STL, concept embeddings are learnt in an unsupervised manner through a cross-attention mechanism. The network makes class predictions through these concept embeddings, allowing for a meaningful explanation to be naturally extracted for each input. Our preliminary experiments with simple CPS-based datasets show that our model is able to achieve great classification performance while ensuring local interpretability.
@@ 3 ECATS architecture
Here we introduce ECATS (Explainable-by-design Concept-based Anomaly detection model for Time Series), the first interpretable neuro-symbolic concept-based model for trajectories. The goal of our proposed methodology is to discriminate between normal and anomalous signals while simultaneously providing a contextual explanation for the detected outcome. The form of the explanation in ECATS is a disjunction of interpretable Signal Temporal Logic(STL) formulae, which function as concepts in our model, and whose semantics influence the labeling of the input trajectories. As mentioned in Section 2.0.2, STL is a concise yet expressive language suited for describing the characteristics of time series data, making it a valuable candidate to serve as a concept language in this field. In fact, small STL formulae describe simple patterns in a signal, and are easy to understand for a human (possibly once converted into a natural language description).
We now describe at a high level how the model works. We provide Figure 1 as a reference to graphically follow the discussion. Our model takes as input a signal $\xi$ and, through a cross-attention mechanism (Subsection 3.1) between a kernel embedding of the STL-based concepts (computed as per Equation 1) and the input trajectory, creating an efficient and detailed representation of their relationship. The representation is then passed to an Multi-Layer Perceptron (MLP) classifier network, which outputs its prediction $y$ indicating the class of the input trajectory. Our model is then able to provide not only excellent classification accuracy (Section 4), but also transparent, easy-to-read explanations made of simple STL formulae (Subsection 3.2).
@@ 4.0.1 Setup
We evaluate ECATS on a commonly-used CPS-based dataset, a train cruise control dataset, often used as learning benchmark [22, train2, train3]. The dataset, balanced between classes, collects $200$ trajectories of $1$ dimension and has a clear-cut condition to classify between regular and anomalous trajectories, except for seven outliers. See Figure 2 for a visualisation of the trajectories of the dataset, with the anomalies highlighted in red.
Figure 2: Train cruise control trajectories. Blue trajectories are regular, red trajectories are anomalous.
@@ 4.0.2 Evaluation
To assess the classification performance of our model, we run the tests using $5$ different initialisation seeds.
For the train cruise control dataset, the classification accuracy averages at $96.5\%$, with a standard deviation of $2.55$, not reaching perfection most likely due to the outliers. Following, some example outputs from ECATS, starting from a local explanation.
Consider the post-processed explanation given for a signal $\xi$ correctly classified as regular, which can be visualised in Figure 3:
@@
We further analyse this result by reporting the plot of the robustness values for the considered signal versus all the signals of the opposite class (Figure 4).
Figure 3: The regular trajectory in the train cruise control dataset considered as example.
Figure 4: Robustness values for the explanation of the regular trajectory reported as example.
As per Figure 4, the provided formula does an excellent job at separating the signal belonging to the regular class from the ones that are anomalous. Coincidentally the cluster of values regarding the anomalous class has opposite sign with respect to the robustness on the considered regular signal, therefore the formula has no need to be shifted to increase human readability.
$\mathcal{G}_{[0,24]}\ \xi\leq 37.3$
$\neg\,\mathcal{G}_{[0,24]}\ \xi\leq 28.7$
See Appendix D for more details on the simplification of formulae.
Figure 5: The anomalous trajectory in the train cruise control dataset considered as example.
Lastly, here is an example of post-processed global explanations given by ECATS:
@@
In Figure 7, the robustness values computed for all trajectories for each formula provided as explanation, where each point in the plot represents the value for one trajectory. Except for some points which are most likely the aforementioned outliers, the class separation is linear.
Figure 7: Robustness of the ECATS’s global explanation for regular (top) and anomalous (bottom) trajectories’ class.
For training details, further discussion of results on the train cruise control dataset, and some preliminary results on a 2-variable dataset, we refer to Appendix E.
@@ 5 Conclusions
{credits}
5.0.1 Acknowledgements
This study was carried out within the PNRR research activities of the consortium iNEST (Interconnected North-Est Innovation Ecosystem) funded by the European Union Next-GenerationEU (Piano Nazionale di Ripresa e Resilienza (PNRR) – Missione 4 Componente 2, Investimento 1.5 – D.D. 1058 23/06/2022, ECS_00000043) and the MUR PRIN project 20228FT78M DREAM (modular software design to reduce uncertainty in ethics-based cyber-physical systems). This manuscript reflects only the Authors’ views and opinions, neither the European Union nor the European Commission can be considered responsible for them.
@@ Appendix A Signal Temporal Logic
Signal Temporal Logic (STL) [19] is a formal specification language to express temporal properties over real-value trajectories with a dense time interval. With STL, one can state in a rich but compact way properties of signals that vary over time. For example, in STL we can express properties like "For the next 3 days the highest temperature will be below 75 degree and the lowest temperature will be above 60 degree" or "There must exist a time point in the next 10 time units, at which the velocity of the car will be less than 2". We now present its syntax and semantics.
@@
Referring to the example mentioned above, we can translate them from plain English to STL as follows: "For the next 3 days the highest temperature will be below 75 degree and the lowest temperature will be above 60 degree" is $G_{[0,3]}(T<75\wedge T>60)$, while "There must exist a time point in the next 10 time units, at which the velocity of the car will be less than 2" is $F_{[0,10]}(v<2)$.
@@ Appendix C Selecting the concepts
Once all the templates have been generated, we instantiate each of them on a set of parameter configurations $\mathcal{P}=\{p_{0},\ldots p_{|\mathcal{P}|}\}$ to be filtered by Signature-based Optimisation as defined in [20]. Denoting as $\varphi(p)$ the application of the parameter set $p$ to the template formula $\varphi$, and given a set of trajectories $\hat{\mathcal{T}}=\{\xi_{0},\ldots,\xi_{|\hat{\mathcal{T}}|}\}$ (with $p\in\mathcal{P}$, and $\hat{\mathcal{T}}\subset\mathcal{T}$), it consists in building a matrix $S\in\mathbb{R}^{|\mathcal{P}|\times|\hat{\mathcal{T}}|}$, whose rows are called the signature of the formula, s.t. $S(i,j)=\rho(\varphi(p_{i}),\xi_{j})$, i.e. the cell $(i,j)$ of such matrix holds the robustness of the $i$-th instantiation of $\varphi$ on the $j$-th trajectory. In such matrix, the distance between rows can be considered as a proxy for semantic similarity among different parameterisations of the same template, hence a selection on formulae can be done by keeping only those whose signature has a distance higher than a certain threshold $\tau\in\mathbb{R}_{\geq 0}$ from formulae already selected. This helps in removing redundancy in the set of concepts. In our case, we use signature-based optimisation with cosine distance and threshold $\tau=0.9$ and, for interpretability, we keep $M$ fairly small, namely $M=3$. Hence, with such procedure, we select as concepts only a subset of all the possible instantiations of each template, following semantic diversity as criterium. Once signature-based optimization is done on all templates, we have the full set of concepts, from which we obtain concept vectors by computing their kernel embeddings.
@@ Appendix E Training setup and results
Let’s see some other local explanations provided by ECATS, starting from one provided for a regular trajectory.
As earlier, consider the post-processed explanation given for a signal $\xi$ correctly classified as regular, which can be visualised in Figure 8:
$\mathcal{G}_{[0,24]}\mathcal{F}_{[0,12]}\,\xi\leq 37.26$
Keeping Figure 2 at hand, this formula doesn’t seem to describe any particular class since in the considered time interval most trajectories are $\leq 37.26$. However, by plotting the robustness (Figure 9) we can see that the formula does discriminate between classes (except for the usual outliers). To aid interpretability, we can shift the threshold, obtaining
$\mathcal{G}_{[0,24]}\mathcal{F}_{[0,12]}\,\xi\leq 27.44$
Now the formula does indeed discriminate between this regular and the other anomalous trajectories while characterising the studied signal.
Figure 8: The regular trajectory in the train cruise control dataset considered as example in appendix.
Figure 9: Robustness values for the explanation of the regular trajectory considered in appendix.
To further prove the capabilities of ECATS, another example of explanation of an anomalous trajectory, reported in Figure 10. The provided explanation for this classification is the following:
$\mathcal{G}_{[0,36]}\,\xi\geq 37.24$
As can be seen from the robustness values Figure 11, this does discriminate between the considered anomaly and the general regular trajectory. In fact, while at the beginning no trajectories satisfy the condition, towards the last points of the interval the considered anomaly does. This proves the condition to be in fact a good discriminant.
Figure 10: The anomalous trajectory in the train cruise control dataset considered in appendix.
Figure 11: Robustness values for the explanation of the anomalous trajectory reported as example in appendix.
E.1 Preliminary results with two variables
As part of this ongoing research, ECATS has been tentatively applied to a second dataset. This too commonly used [22, nenzi2018robust, bombara], it consists of a maritime surveillance scenario, presenting a more challenging task due to the absence of distinct criteria for distinguishing between regular and anomalous trajectories. The dataset collects $2000$ trajectories, balanced between classes, of $2$ dimensions. Trajectories consist of the longitudinal and latitudinal coordinates ($x_{0},x_{1}$) of vessels transiting over a harbour, with anomalous trajectories corresponding to vessels involved in illegal activities, and regular trajectories corresponding to vessels following the expected route. See Figure 12 for a visualisation of the trajectories of the dataset, with the anomalies highlighted in red. For better understanding of the explanations, we provide both a view of the variables as coordinates of points on a 2D plane (12(a)) and of the value of the single variables related to time (12(b)).
(a) View as plane coordinates.
(b) View with reference to time.
Figure 12: Maritime surveillance trajectories. Blue trajectories are regular, red trajectories are anomalous.
Note that the following results are preliminary and the model is still to be tweaked to adapt to more complex problems.
For the maritime dataset, ECATS reaches $100\%$ classification accuracy in all runs. Following, some example outputs for local explanations, starting from a regular trajectory. Consider the post-processed explanation given for a signal $\xi=\{x_{0},x_{1}\}$ correctly classified, which can be visualised in Figure 13. The provided explanation is the following:
$\mathcal{F}_{21,51]}x_{1}\geq 28.38\ \land\ \mathcal{G}_{[35,\infty]}x_{0}\leq 18.8$
This conjunction of formulas does correctly identify the trajectory. The left one is false for half of the anomalous trajectories, while the right one is false for the other half.
Figure 13: The regular trajectory in the maritime dataset considered as example.
Figure 14: Robustness values for the explanation of the regular trajectory.